OpenStudy (anonymous):

What is the Bandwidth of sinc(100*pi*t)

OpenStudy (anonymous):

I used the Fourier transform to get : \[\frac{ 1 }{ 100\pi }rect(\frac{ f }{ 100\pi })\] if you graph this you get bandwidth = 50*pi Hz but the answer is apparently 50 Hz can someone please explain this

OpenStudy (kenljw):

The original function is in radian frequency, must divide by 2pi to convert to HZ